Circular No. 5731

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 J. Ripero, Madrid, Spain, reports that F. Garcia, Lugo, Spain, dis-

covered a possible supernova on Mar. 28 as much as 5′ southwest of the nucleus of NGC 3031 = M81. The object was also recorded in an ST-4 CCD image obtained by D. Rodriguez as some 30′′ northeast of a mag 14 foreground star. Magnitude estimates: Mar. 26.9 UT, [14.0 (Garcia and P. Pujol); 28.86, 12.0 (Garcia, visual), 29.1, 11.8 (Rodriguez, CCD unfiltered); 29.88, 11.3 (Pujol); 29.88, 11.0 (Rodriguez).

A. V. Filippenko, University of California at Berkeley, reports: “A CCD image of M81 obtained on Mar. 30.1 UT by R. R. Treffers and Y. Paik (also of Berkeley) with the 0.8-m reflector at Leuschner Observatory confirms the presence of a new stellar object ∼ 45′′ west and 160′′ south of the nucleus. The visual magnitude is ∼ 11. Inspection of CCD spectra (range 356-731 nm) obtained on Mar. 30.3 by M. Davis and D. Schlegel (also of Berkeley) with the Lick 3-m Shane reflector reveals that the object is indeed a supernova. The continuum is very blue and remarkably featureless. The only clear absorption lines are narrow Na I D and Ca II H + K, undoubtedly of interstellar origin. It is probable, but not yet certain, that the object is a type II supernova observed only a few days after the explosion. Note, however, that the type Ia SN 1991T exhibited a relatively featureless spectrum well before maximum brightness (Filippenko et al. 1992, Ap.J. 384, L15). Depending on its spectral type, distance and extinction, SN 1993J may reach eighth magnitude during the next two weeks. Aside from SN 1987A in the LMC, it is therefore the brightest supernova since SN 1972E in NGC 5253. Further observations throughout the electromagnetic spectrum are urged.”

F. D. A. Hartwick, D. D. Balam, D. Zurek and R. M. Robb, Climenhaga Observatory, University of Victoria, provide the following precise position for the supernova, measured by Balam from a CCD image obtained with the 0.5-m reflector on Mar. 30.25 UT: α = 9^{h}51^{m}19^{s}.27, δ = +69^{o}15′25′′.7 (equinox 1950.0). Photometry yields V = 10.2 ± 0.1.

P. Garnavich and B. A. Hong, Dominion Astrophysical Observatory, communicate: “Spectra (range 550-690 nm, resolution 0.6 nm) of the supernova were taken with the DAO 1.8-m telescope on Mar. 30.25 UT. The spectrum shows a strong, flat continuum with weak Hα and He I (587.5 nm) features, consistent with that of a type II supernova.”

1993 March 30 Brian G. Marsden